40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Tue Dec 6 02:57:44 2011, kiwamu, Update, Green Locking, 1st trial of handing off YarmALS.png
    Reply  Wed Dec 7 02:55:38 2011, kiwamu, Update, Green Locking, locking activity tonight 
Message ID: 6080     Entry time: Wed Dec 7 02:55:38 2011     In reply to: 6076
Author: kiwamu 
Type: Update 
Category: Green Locking 
Subject: locking activity tonight 

No real progress.

Probably I spent a bit too much time realigning the beat-note optical path.

 

(what I did)

 - Switched on a power supply which was supposed to give +/- 15V for the broadband beat-note PD.
   The power supply had been somehow turned off.
 - Realigned the beat-note path. When we installed the new EOM mount today, we moved some of the green steering mirrors to make a space.
   So we had to realign the downstream of the beat-note path. After the realignment the DC output of the PD was about 120 mV and the signal level of the beat-note was at -20 dBm.
 - Took noise spectrum of the beat-note with the arm cavity locked by the IR-PDH
    The noise curve was almost the same as before (i.e. unknown high frequency white noise above 20 Hz and some low frequency noise which has structures at 1 and 3 Hz).
-  Closed the ALS loop with the coarse sensor. But I was too lazy to go further more. 

Quote from #6076
Tomorrow I will try :
  (1) Using the fine sensor.
  (2) Noise budgeting with the fine sensor.

ELOG V3.1.3-